OpenStudy (hhhhhjauhsd):

WILL FAN AND MEDAL Maria plans to use $18,000 to open a savings account with an annual interest rate of 7%. How much more interest will she earn over 6 years if she chooses an account that compounds interest quarterly instead of annually?

OpenStudy (hhhhhjauhsd):

@pooja195

OpenStudy (hhhhhjauhsd):

@Preetha

satellite73 (satellite73):

compare \[18,000\times (1.07)^6\] with \[18,000\times (1+\frac{.07}{4})^{4\times 6}\]

OpenStudy (hhhhhjauhsd):

439,560?

OpenStudy (hhhhhjauhsd):

@satellite73

satellite73 (satellite73):

idk i didn't do it

satellite73 (satellite73):

here is the first one http://www.wolframalpha.com/input/?i=18,000*%281.07%29^6

satellite73 (satellite73):

here is the second one http://www.wolframalpha.com/input/?i=18,000*%281%2B.07%2F4%2924